OpenStudy (anonymous):

A student solved this problem and said the answer is 1 and 1/6 miles. Jeff ran on two different days. The first day he ran 1/3 mile and the second day he ran 5/6 mile. What is the total distance Jeff ran? Is the student's answer reasonable? A. Yes, the answer is reasonable. B. No, the answer is not reasonable. It should be about 2 miles. C. No, the answer is not reasonable. It should be about 1/2 mile. D. No, the answer is not reasonable. It should be about 3 miles.

OpenStudy (camerondoherty):

Change 1/3 to 2/6 first

OpenStudy (camerondoherty):

then add

OpenStudy (camerondoherty):

2/6 +5/6=

OpenStudy (thomas5267):

Firstly, define "reasonable" lol. Secondly, \(\dfrac{1}{3}=\dfrac{1\times 2}{3\times2}=\dfrac{2}{6}\).

OpenStudy (camerondoherty):

1 and 1/6

OpenStudy (camerondoherty):

so it is reasonable

OpenStudy (anonymous):

Thank you! My brain is having multiple brain farts o.O